# Ion Coulomb crystals: an exotic form of condensed matter

### v1

@@ II.3.3 Normal modes of three dimensional crystals
Here $Q^{m}_{l}$ and $P^{m}_{l}$ are associated Legendre functions, the coordinate $\xi_{1}$ is a generalized distance coordinate (similar to spherical radius), $\xi_{2}$ is a generalized latitude (similar to the polar angle $\theta$), $\phi$ is the usual azimuthal angle, $\omega$ is the mode frequency, and $d^{2}\equiv Z_{p}^{2}-R_{p}^{2}$, where $2Z_{p}$ is the axial extent and $2R_{p}$ is the diameter of the spheroid. The specification of the mode potential in the interior of the ion crystal is more complicated.
@@
Figure 38: Normal mode frequencies $\omega$ in the rotating frame of the ion crystal vs rotation frequency $\omega_{r}$ for $\omega_{z}/\Omega=0.151$ and different values of $(l,m)$. Frequencies $\omega<0$ are allowed and obtained in the diagrams by reflecting through the point $(\omega_{r},\omega)=(\Omega/2,0)$. For $m\neq 0$ the sign of $\omega$ determines the direction of the azimuthal motion of the mode potential. Also shown for comparison are the vortex frequency $\Omega_{v}(\omega_{r})=\Omega-2\omega_{r}$ (dotted line), the plasma frequency $\omega_{\mathrm{p}}(\omega_{r})$ (see Eq. (4)) (dot-dashed curve), and the upper hybrid frequency $\Omega_{u}(\omega_{r})=(\omega_{\mathrm{p}}^{2}+\Omega_{v}^{2})^{1/2}$ (solid curve). All frequencies are expressed in units of the cyclotron frequency $\Omega$. (a) $(l,m)=(2,1)$, filled circles; $(l,m)=(5,1)$, open circles. (b) $(l,m)=(2,0)$, filled circles; $(l,m)=(4,0)$, open circles. From 28.
@@
Without going into further details we discuss some example calculations of $m=1$ and $m=0$ modes in Figs. 38(a) and (b), respectively. Both figures show the calculated mode frequency $\omega$ normalized to the ion cyclotron frequency $\Omega$ as a function of the ion crystal rotation $\omega_{r}$ in a Penning tap where $\omega_{z}/\Omega=0.151$, typical of NIST experimental work. (Note that in Secs. I.2.2 and II.2 we used $\Omega_{c}$ to denote the ion cyclotron frequency.) The mode frequencies are plotted in the rotating frame of the ion crystal. For $\omega_{r}/\Omega\ll 1$ (or $\omega_{r}/\Omega\sim 1$) the mode frequencies separate into three different bands. These are the 3D analogues of the ${\mathbf{E}}\times{\mathbf{B}}$, drumhead, and cyclotron modes discussed in the planar geometries section (Sec. II.2.3). A difference with respect to the planar modes is that the ${\mathbf{E}}\times{\mathbf{B}}$ (and to a much lesser extent the cyclotron modes) now acquire a component of motion that is parallel to the magnetic field [97, 138].
@@
As $\omega_{r}$ increases from $\omega_{r}/\Omega\ll 1$, the separation of the modes into three bands becomes less distinct. At $\omega_{r}/\Omega=0.5$ the effective magnetic field (given by the so-called vortex frequency) in the rotating frame of the ion crystal goes to zero. At this point the dynamics of the ion crystal is unmagnetized and the calculated mode frequency and potential are relevant for a spheroidal ion crystal in an rf trap with the same shape or aspect ratio as the Penning trap ion crystal. In this way modes of spheroidal crystals in rf traps can be catalogued by the same values of $(l,m)$ as in the Penning trap [67]. This discussion assumes the validity of the pseudopotential approximation. For a discussion about the validity of this approximation and a mathematical analysis of modes that includes the time-dependent dynamics of rf Paul traps, see Refs. [181, 182].
@@
Experimentally, the normal modes of 3D trapped ion crystals have been excited by applying time-dependent potentials to the trap electrodes [215, 67] and through the application of radiation pressure forces [180]. Typically, the ion motion produced by exciting the mode is detected through Doppler shifts. In Ref [67] several $(l,m=0)$ mode frequencies of spheroidal ion crystals stored in a rf trap were measured through rf excitation and detection of the increased Doppler shifts through the resulting modification of the coupling of the ion electronic states to an optical cavity. Excellent agreement (better than 1%) with the predictions of the cold, spheroidal mode theory of Dubin [93] was observed. In Penning traps, rf excitation and detection of the change in the ion fluorescence from Doppler shifts has been used to measure modes $(l,m)$ with $l$ as large as 9 and $m$ as large as 2 [140, 214]. We note that sideband spectroscopy and detection of ion motion through the application of spin-dependent optical dipole forces, techniques used to detect modes with linear ion crystal strings and 2D ion crystals, will also work for detecting normal modes in 3D crystals.
@@
Figure 39: Sideview image of the velocity profile of an $(l,m=0)$ mode excited in a spheroidal, laser-cooled cloud of 9Be+ ions in a Penning trap. Here $\omega_{r}/(2\pi)=1.00$ MHz, $\omega_{z}/(2\pi)=1.13$ MHz, and $\Omega/(2\pi)=7.608$ MHz (cyclotron frequency). The magnetic field and axial laser beam point up. The ion cloud spheroid dimensions are $2Z_{p}=0.76$ $m$ and $2R_{p}=0.24$ $m$, with the density $n_{0}=2.70\times 10^{9}$ cm-3. At each point of the image, the measured ion resonance fluorescence is fitted to an amplitude and phase. (a) The left panel compares the measured amplitude of a (2,0) mode excited at $\omega_{2,0}/(2\pi)=1.656$ MHz with the predictions of linear theory. The right side compares the measured phase with theory. (b) The left panel compares the measured amplitude of a (9,0) mode excited at $\omega_{9,0}/(2\pi)=2.952$ MHz with the predictions of linear theory. The right side compares the measured phase with theory. From 215.
@@
In addition to measuring mode frequencies, an image of the excited mode eigenfunction can be obtained through an image of the laser-induced ion fluorescence from a trapped ion crystal. Figure 39 shows experimentally measured images of $(2,0)$ and $(9,0)$ modes excited in cold, spheroidal clouds of 9Be+ ions in a Penning trap. Here a laser beam is directed parallel to the magnetic field axis (the vertical direction in Fig. 39; see Fig. 47 for an experimental schematic) with a frequency tuned below resonance with a cycling transition of 9Be+. Ions moving with a velocity in the direction opposite to the laser k vector fluoresce more strongly than ions moving in the same direction as the laser k vector. By recording side-view images of the ion crystal with a photon-counting imaging camera that measures the spatial and temporal coordinates of each detected photon, it is possible to construct side-view images of the excited mode velocity profile that are phase synchronous with the external rf drive used to excite the mode. Figure 39 shows images of excited $(2,0)$ and $(9,0)$ mode eigenfunctions that are in good agreement with theory.
III Out of equilibrium dynamics of ion Coulomb crystals
@@
Understanding how strongly-correlated many-body systems reach equilibrium is central to statistical physics and is essential for applications to quantum technologies. By offering a controllable environment, ion Coulomb crystals are a prominent platform for simulating models that allow verification of equilibration and thermalization hypotheses. Beyond engineering, the intrinsic physical properties make the system peculiar by itself. In fact, the unscreened Coulomb interactions make ion Coulomb crystals a unique laboratory for studying the out-of-equilibrium dynamics of long-range interacting systems, which can clarify hypotheses and conjectures formulated for cosmological and plasma systems [45, 72]. In this section we review experimental and theoretical work on out-of-equilibrium dynamics in ion Coulomb crystals. We focus in particular on transport and relaxation of Coulomb crystals after a perturbation, such as a slow or sudden change (quench) of the trapping parameters or of the temperature. One relevant aspect is the formation, stability and decay of defects and dislocations, which in ion crystals can be imaged both in real as well as in Fourier space. As the concept of out-of-equilibrium is central to this section, we start by defining what can be understood as equilibrium in a trapping environment, and when the laser-cooled crystal can be said to be in a thermal state.
@@ III.1 Equilibrium and temperature
In the Penning trap (see Sec. I.2), the trapping forces break time reversal symmetry and confinement is achieved by setting the ions in rotation. Collisions between the ions drive the system into a state where the ion ensemble rotates rigidly (without shear) about the symmetry axis of the trap, which is parallel to direction of the magnetic field. Rigid body rotation has been observed in the laboratory [89, 34]. In the frame co-rotating with the ions, the system can evolve into a thermal equilibrium state characterized by a global temperature [98].
@@
On longer time scales, anharmonicities can redistribute the energy between the modes leading to a thermal state [226, 224]. At first glance, because radio frequency (rf) and Penning ion traps can routinely confine ensembles of charges of the same sign for many hours to days, the achievement of global thermal equilibrium states appears feasible. However, depending on the details of the trap parameters, the time to equilibrate to a global thermal state can exceed even these time scales. Consider, for example, a long linear chain of ions. Here the confinement in the transverse direction, characterized by frequency $\nu_{t}$ (see Eq. (27)) is much stronger than the confinement in the longitudinal direction, characterized by frequency $\nu$. This results in normal mode frequencies for motion transverse to the trap axis that are of much higher frequency than the normal mode frequencies that describe motion parallel to the trap axis. In this case it has been shown theoretically that the rate of equilibration between the transverse and parallel motional degrees of freedom can be exponentially small in the ratio $\nu_{t}/\nu$, leading to equilibration time periods that exceed many months for cold ion chains with $N>10$ [56]. Physically, this result can be understood through energy conservation arguments. For large ratios $\nu_{t}/\nu$ the coupling between the transverse (optical) and the longitudinal (acoustic) bands is a weak perturbation. In terms of vibrational quanta an interband transition requires that several low-frequency acoustic phonons are annihilated to create a single high-frequency optical phonon. Such multi-phonon processes arise as high-order terms in the Coulomb interaction and have small effective rates.
@@
Long equilibration rates are also anticipated with single plane crystals with large ion number $N$, which require a trap asymmetry parameter $\beta\ll 1$ (see Eq. (55)). Lastly, in Penning traps, the strong magnetic field splits the normal modes that describe in-plane motion into high frequency cyclotron modes and low frequency ${\mathbf{E}}\times{\mathbf{B}}$ modes (see Fig. 25). Numerical studies show that the equilibration of the cyclotron and ${\mathbf{E}}\times{\mathbf{B}}$ branches is exponentially weak in the ratio of the frequencies of the two branches [309].
@@
In summary, although ion Coulomb crystals can be confined for long times, this might not be sufficient to observe equilibration between different bands, expecially when these are separated by relatively large energy gaps. Equilibration processes can be experimentally accelerated by various strategies, such as fine tuning of the trap parameters. As a simple example, for a two-ion crystal aligned along the axis of a linear rf trap, the axial stretch mode frequency can be tuned to be approximately equal to twice the frequency of the radial rocking mode. In this case the intrinsic non-linear Coulomb force between the ions leads to a strong non-linear coupling between these two modes, as illustrated in Fig. 40 [79]. With a linear three-ion crystal it has furthermore been possible to tune and exploit the resonant coupling of three normal modes  [201]. In Penning traps, numerical studies show that the coupling between in-plane ${\mathbf{E}}\times{\mathbf{B}}$ modes and the out-of-plane drumhead modes for single plane crystals can be strongly enhanced by tuning trap parameters so that their respective bandwidths overlap [161].
@@
Figure 40: Parametric coupling between an axial ($a$) and a radial ($r$) normal mode of a two-ion system. The mode frequencies are tuned such that $\omega_{a}=2\omega_{r}$. The red data points indicate the time evolution of the excitation of the axial (solid dots) and the radial (open circles) mode when the system initially is prepared with two quanta of excitation in the radial mode. In this case, a coherent exchange of energy is observed between the two modes. In case the radial mode is initially prepared with only one quantum of excitation (blue data points), no coherent exchange is observed due to the lack of energy conservation. The error bars show one standard deviation statistical uncertainty. From 79.
We conclude by discussing temperature measurement in rf-traps. Temperature is determined by extracting the kinetic energy associated with the rf-driven coherent micromotion, which is typically several orders of magnitude larger than the thermal motion. This can be done, e.g., in Molecular Dynamics (MD) simulations, by recording the change in the individual ion velocities at the same phase of the rf fields over time [279]. By such MD simulations, one can also obtain knowledge of the heating of the Coulomb crystals due to the rf drive [26, 279, 310, 232]. In experiments with 3D Coulomb crystals, where the kinetic energies of the individual ions (or equivalently the various normal mode excitations) are difficult to measure, comparison between the experimentally measured fluorescence images and the ones generated from MD simulations are often a practical way to estimate the temperature [241, 239, 174].

### v2

@@ II.3.3 Normal modes of three-dimensional crystals
Here, $Q^{m}_{l}$ and $P^{m}_{l}$ are associated Legendre functions, the coordinate $\xi_{1}$ is a generalized distance coordinate (similar to spherical radius), $\xi_{2}$ is a generalized latitude (similar to the polar angle $\theta$), $\phi$ is the usual azimuthal angle, $\omega$ is the mode frequency, and $d^{2}\equiv Z_{p}^{2}-R_{p}^{2}$, where $2Z_{p}$ is the axial extent and $2R_{p}$ is the diameter of the spheroid. The specification of the mode potential in the interior of the ion crystal is more complicated.
@@
Figure 38: Normal mode frequencies $\omega$ in the rotating frame of the ion crystal vs rotation frequency $\omega_{r}$ for $\omega_{z}/\Omega=0.151$ and different values of $(l,m)$, where $\Omega\equiv\Omega_{c}$ is the cyclotron frequency. Frequencies $\omega<0$ are allowed and obtained in the diagrams by reflecting through the point $(\omega_{r},\omega)=(\Omega/2,0)$. For $m\neq 0$ the sign of $\omega$ determines the direction of the azimuthal motion of the mode potential. Also shown for comparison are the vortex frequency $\Omega_{v}(\omega_{r})=\Omega-2\omega_{r}$ (dotted line), the plasma frequency $\omega_{\mathrm{p}}(\omega_{r})$ (see Eq. (4)) (dot-dashed curve), and the upper hybrid frequency $\Omega_{u}(\omega_{r})=(\omega_{\mathrm{p}}^{2}+\Omega_{v}^{2})^{1/2}$ (solid curve). (a) $(l,m)=(2,1)$, filled circles; $(l,m)=(5,1)$, open circles. (b) $(l,m)=(2,0)$, filled circles; $(l,m)=(4,0)$, open circles. From 31.
@@
We discuss some example calculations of $m=1$ and $m=0$ modes in Figs. 38(a) and (b), respectively. Both figures show the calculated mode frequency $\omega$ normalized to the ion cyclotron frequency $\Omega$ ($\Omega\equiv\Omega_{c}$ in Secs. I.2.2 and II.2) as a function of the ion crystal rotation $\omega_{r}$ in a Penning trap with axial frequency $\omega_{z}/\Omega=0.151$. The mode frequencies are plotted in the rotating frame of the ion crystal. For $\omega_{r}\sim 0$ (or $\omega_{r}\sim\Omega$) the ion crystal has a lenticular shape (see Fig. 7) and the mode frequencies separate into three different bands. These bands are the 3D analogues of the ${\mathbf{E}}\times{\mathbf{B}}$, drumhead, and cyclotron modes discussed in the planar geometries section II.2.3. A difference with respect to single-plane modes is that the ${\mathbf{E}}\times{\mathbf{B}}$ (and to a much lesser extent the cyclotron modes) now acquire a component of motion that is parallel to the magnetic field [103, 150].
@@
As $\omega_{r}/\Omega$ increases from zero, the separation of the modes into three bands becomes less distinct. At $\omega_{r}/\Omega=0.5$ the effective magnetic field in the rotating frame of the ion crystal goes to zero. At this special point the equations of motion of the normal modes are equivalent to the ones of a spheroidal ion crystal with the same aspect ratio in an rf trap in the psuedopotential approximation, and the modes can be classified with the integers $(l,m)$ of Eq. (65[72]. For a discussion about the limits of validity of this equivalence and a mathematical analysis of modes that includes the time-dependent dynamics of rf Paul trap potentials, see Refs. [198, 199].
@@
Experimentally, the normal modes of 3D trapped ion crystals have been excited by applying time-dependent potentials to the trap electrodes [239, 72] or by means of radiation pressure [196]. Typically, the ion motion produced by exciting the mode induces Doppler shifts that can be used to detect the mode. In Ref [72] several $(l,m=0)$ modes of spheroidal ion crystals stored in a rf trap were excited through rf excitation and the resulting Doppler shifts detected through modification of the collective coupling of the ion crystal to an optical cavity. Excellent agreement (better than 1%) was observed with the mode frequency predictions of the cold, spheroidal theory of Dubin [99]. In Penning traps, detection of the Doppler shifts produced by an excited mode through changes in the ion fluorescence (so-called Doppler velocimetry) was used to measure the frequencies of modes $(l,m)$ with $l$ as large as 9 and $m$ as large as 2 [152, 238].
@@
Figure 39: Sideview image of the velocity profile of an $(l,m=0)$ mode excited in a spheroidal, laser-cooled cloud of 9Be+ ions in a Penning trap. Here, $\omega_{r}/(2\pi)=1.00$ MHz, $\omega_{z}/(2\pi)=1.13$ MHz, and $\Omega/(2\pi)=7.608$ MHz, where here $\Omega\equiv\Omega_{c}$ is the cyclotron frequency. The magnetic field and axial laser beam point up. The ion cloud spheroid dimensions are $2Z_{p}=0.76$ $m$ and $2R_{p}=0.24$ $m$, with the density $n_{0}=2.70\times 10^{9}$ cm-3. At each point of the image, the measured ion resonance fluorescence is fitted to an amplitude and phase. (a) The left panel compares the measured amplitude of a (2,0) mode excited at $\omega_{2,0}/(2\pi)=1.656$ MHz with the predictions of linear theory. The right panel compares the measured phase with theory. (b) The left panel compares the measured amplitude of a (9,0) mode excited at $\omega_{9,0}/(2\pi)=2.952$ MHz with the predictions of linear theory. The right panel compares the measured phase with theory. From 239.
@@
Images of the excited mode eigenfunction of a trapped ion crystal can be obtained through the laser-induced ion fluorescence. Figure 39 shows experimentally measured images of $(2,0)$ and $(9,0)$ modes excited in cold, spheroidal clouds of 9Be+ ions in a Penning trap. Here, a laser beam is directed parallel to the magnetic field axis (the vertical direction in Fig. 39) with a frequency tuned below resonance with a cycling transition of 9Be+, see Fig. 47 for the experimental schematic. Ions moving with a velocity in the direction opposite to the laser wave vector k fluoresce more strongly than ions moving in the same direction as k. By recording side-view images of the ion crystal with a photon-counting imaging camera that measures the spatial and temporal coordinates of each detected photon, it is possible to construct side-view images of the excited mode velocity profile that are phase synchronous with the external radio frequency drive used to excite the mode. Comparison of the experimental images with the theoretical predictions in Fig. 39 shows that experiment and theory are in good agreement.
III Out-of-equilibrium dynamics of ion Coulomb crystals
@@
Understanding how strongly-correlated many-body systems reach equilibrium is a central objective of statistical physics and is essential for applications to quantum technologies. By offering a controllable environment, ion Coulomb crystals are a prominent platform for simulating models that allow verification of equilibration and thermalization hypotheses. The intrinsic physical properties make the system peculiar by itself. In fact, the unscreened Coulomb interactions make ion Coulomb crystals a unique laboratory for studying the out-of-equilibrium dynamics of long-range interacting systems, which can clarify hypotheses and conjectures formulated for cosmological and plasma systems [49, 77]. In this section we review experimental and theoretical work on out-of-equilibrium dynamics in ion Coulomb crystals. We focus in particular on transport and relaxation of ion Coulomb crystals after a perturbation from equilibrium, such as a slow or sudden change (quench) of the trapping parameters or of the temperature. One relevant aspect is the formation, meta-stability and decay of defects and dislocations, which in ion crystals can be imaged both in real as well as in Fourier space. As the concept of out-of-equilibrium is central to this section, we start by defining what can be understood as equilibrium in a trapping environment, and when the laser-cooled crystal can be said to be in a thermal state.
@@ III.1 Equilibrium and temperature
In the Penning trap (see Sec. I.2), the trapping forces break time reversal symmetry and confinement is achieved by setting the ions in rotation. Collisions between the ions drive the system into a state where the ion ensemble rotates rigidly (without shear) about the symmetry axis of the trap, which is parallel to the direction of the magnetic field. Rigid body rotation has been observed in the laboratory [37, 95]. In the frame co-rotating with the ions, the system can evolve into a thermal equilibrium state characterized by a global temperature [105].
@@
On longer time scales, anharmonicities can redistribute the energy between the modes leading to a thermal state [252, 250]. At first glance, because radio frequency (rf) and Penning traps can routinely confine ensembles of charges of the same sign for many hours to days, the achievement of global thermal equilibrium states appears feasible. However, depending on the details of the trap parameters, the time to equilibrate to a global thermal state can exceed even these time scales. Consider, for example, a long linear chain of ions. Here, the confinement in the transverse direction, characterized by frequency $\nu_{t}$ (see Eq. (27)) is much stronger than the confinement in the longitudinal direction, characterized by frequency $\nu$. This results in normal mode frequencies for motion transverse to the trap axis that are much higher than the normal mode frequencies of motion parallel to the trap axis. In this case it has been shown theoretically that the rate of equilibration between the transverse and parallel motional degrees of freedom can be exponentially small in the ratio $\nu_{t}/\nu$, leading to equilibration time periods that exceed many months for cold ion chains with $N>10$ [61]. Physically, this result can be understood through energy conservation arguments. For large ratios $\nu_{t}/\nu$ the coupling between the transverse (optical) and the longitudinal (acoustic) bands is a weak perturbation. In terms of vibrational quanta an interband transition requires that several low-frequency acoustic phonons are annihilated to create a single high-frequency optical phonon. Such multi-phonon processes arise as high-order terms in the Coulomb interaction and have very small effective rates.
@@
Long equilibration rates are also anticipated with single-plane crystals with large ion number $N$, which require a trap asymmetry parameter $\beta\ll 1$ (see Eq. (55)). Lastly, in Penning traps, the strong magnetic field splits the normal modes that describe in-plane motion into high frequency cyclotron modes and low frequency ${\mathbf{E}}\times{\mathbf{B}}$ modes (see Fig. 25). Numerical studies show that the equilibration of the cyclotron and ${\mathbf{E}}\times{\mathbf{B}}$ branches is exponentially weak in the ratio of the frequencies of the two branches [344].
@@
In summary, although ion Coulomb crystals can be confined for long times, this might not be sufficient to observe equilibration between different bands, especially when these are separated by relatively large energy gaps. Equilibration processes can be experimentally accelerated by various strategies, such as fine tuning of the trap parameters. As a simple example, for a two-ion crystal aligned along the axis of a linear rf trap, the axial stretch mode frequency can be tuned to be approximately equal to twice the frequency of the radial rocking mode. In this case the intrinsic non-linear Coulomb force between the ions leads to a strong non-linear coupling between these two modes, as illustrated in Fig. 40 [85]. Furthermore, in a linear crystal of three ions it has been possible to tune and exploit the resonant coupling of three normal modes  [224]. In Penning traps, numerical studies show that the coupling between in-plane ${\mathbf{E}}\times{\mathbf{B}}$ modes and the out-of-plane drumhead modes for single-plane crystals can be strongly enhanced by tuning trap parameters so that their respective bandwidths overlap [174].
@@
Figure 40: (Color online) Parametric coupling between an axial ($a$) and a radial ($r$) normal mode of a two-ion system. The mode frequencies are tuned such that $\omega_{a}=2\omega_{r}$. The data points indicate the time evolution of the excitation of the axial (solid dots) and the radial (open circles) mode when the system is initially prepared with either one or two quanta of excitation in the radial mode. In the latter case (data points fitted by a sinusoidal red solid curve), a coherent exchange of energy is observed between the two modes. In case the radial mode is initially prepared with only one quantum of excitation (data points fitted by solid blue horizontal lines), no coherent exchange is observed and the modes are effectively decoupled. The error bars show one standard deviation statistical uncertainty. From 85.
We conclude by discussing temperature measurement in rf-traps. Temperature is determined by extracting the kinetic energy associated with the rf-driven coherent micromotion, which is typically several orders of magnitude larger than the thermal motion. This can be done, e.g., by recording in molecular dynamics simulations the change in the individual ion velocities at the same phase of the rf fields over time [313]. In this way, one can also gain knowledge of the heating of the Coulomb crystals due to the rf drive [29, 313, 345, 258, 280]. In experiments with 3D Coulomb crystals, where the kinetic energies of the individual ions (or equivalently the various normal mode excitations) are difficult to measure, comparison between the experimentally measured fluorescence images and the ones generated from molecular dynamics simulations are often a practical way to estimate the temperature [268, 266, 189].